# Improving Student Self-Efficacy in Quantum Computing with the Qubit Touchdown Board Game

### v1

@@
Improving Student Self-Efficacy in Quantum Computing with the
@@
Qubit Touchdown is a two-player, competitive board game that was developed to introduce students to quantum computing. A quantum computer is a new kind of computer that is based on the laws of quantum physics, and it can solve certain problems faster than normal computers because it follows a different set of rules. Qubit Touchdown’s game play mirrors the rules of (American) football, with players taking turns moving the football to score the most touchdowns, and no knowledge of quantum computing is needed to play the game. We evaluated the game with 107 public high school students in Precalculus, Advanced Placement (AP) Statistics, and/or AP Physics 1 courses, assessing whether their interest in and self-efficacy toward quantum computing changed as a result of playing the game and learning about its connections to quantum computing. We also assessed whether the game was easy to learn and enjoyable. We found that students’ self-efficacy was improved by 33.4%, and they widely considered the game accessible and fun. Thus, Qubit Touchdown could be an effective resource to introduce students to Quantum Computing and boost their confidence in learning about the field. Free printables of the game are available, and professionally produced copies can be purchased on demand.
@@ III Connections to Quantum Computing
IV Research on Student Interest, Self-Efficacy, and Enjoyment
To evaluate whether Qubit Touchdown can affect students’ interest and self-efficacy in quantum computing, as well as whether the game is easy to learn and fun to play, we conducted a study consisting of 107 public high school students across seven Precalculus, AP Statistics, and/or AP Physics 1 classes over four periods in one school day, which necessitated combining some classes in the library for the activity.
@@ IV Research on Student Interest, Self-Efficacy, and Enjoyment
Table 2: Overall agree and disagree percentages ($p$ and $q$) for the survey results, as well as the corresponding binomial distribution fits ($p^{\prime}$ and $q^{\prime}$) for each survey item.
Question # $p$ $q$ $p^{\prime}$ $q^{\prime}$
Q1 36.4% 26.1% 58.2% 41.8%
Q2 36.4% 29.0% 55.7% 44.3%
Q3 46.7% 25.2% 64.9% 35.1%
Q4 47.7% 15.0% 76.1% 23.9%
Q5 84.1% 3.7% 95.7% 4.3%
Q6 100.0% 0.0% 100.0% 0.0%
Q7 46.7% 22.4% 67.6% 32.4%
Q8 69.2% 8.4% 89.2% 10.8%
Q9 85.0% 1.9% 97.8% 2.2%
Q10 84.1% 3.7% 95.7% 4.3%
Q11 78.5% 5.6% 93.3 % 6.7%
Q12 85.0% 3.7% 95.8% 4.2%
Q13 83.2% 1.9% 97.8% 2.2%
@@
To analyze the data, we treat the Likert scale results as ordinal data and combined agree and strongly agree responses, and disagree and strongly disagree responses, to form overall agree and disagree values, $p$ and $q$, respectively, as shown in Table 2. The overall agree results indicate that students were overall very favorable about their experience of playing Qubit Touchdown. For example, 78.5% of the students reported that playing Qubit Touchdown made quantum computing concepts less intimidating (question #11), and 84.1% of the students reported that they enjoyed playing the game (question #5). Similarly, 83.2% of students recommended that Qubit Touchdown be used as a learning tool for quantum computing (question #13).
To analyze the statistical significance of shifts from pre-game to post-game or post-instruction, we fit the data to a binomial distribution following the method of Redish, Saul, and Steinberg,[21] yielding variables
$p^{\prime}=\frac{p}{p+q}\quad\text{and}\quad q^{\prime}=\frac{q}{p+q}.$ (1)
These values are also shown in Table 2. Standard practice is to consider shifts from pre- to post-test as statistically significant if the difference is less than the 5% probability level, i.e. shifts must be greater than the $2\sigma$-level where $\sigma$ is defined as $\sigma=\sqrt{p^{\prime}q^{\prime}/n}$, where $n$ is the number of students responding to the survey. Since the pre- and post-tests consists of roughly similar questions asked before playing Qubit Touchdown, after playing Qubit Touchdown, and then again after some brief instruction about quantum computing and how it ties to the game, we analyzed shifts between pairs of questions. Table 3 gives the favorable shift (movement towards more agreement) and the level of statistical significance of several question pairs. We found that there were several statistically significant shifts, particularly related to student self-efficacy, i.e., student confidence in their ability to learn quantum computing. Of particular note is the shift in students’ self-confidence in their ability to learn quantum computing as seen in question pairs #2 to #4 and question #2 and question #8. From question #2 to question #4, students’ self-confidence in their ability to learn quantum computing improved by 20.4%, a statistically significant shift of more than $4.2\sigma$. The shift from question #2 to question #8 (pre-game to post-instruction) had a favorable shift of 33.4%, a statistically significant shift of more than $7.0\sigma$. Although we expected students to be more interested in quantum computing after playing Qubit Touchdown, we found that there was not a statistically significant shift in student attitudes, seen in, for example, questions #1 and #3 or questions #3 to #7 (all of which deal with students level of interest in quantum computing). Although there was a small effect when considering the favorable shift from question #1 to question #7 (a 9.4% favorable shift at the $2.1\sigma$ level), we theorize that a greater favorable shift was precluded because the sample included only students in Precalculus, AP Statistics, or AP Physics 1 classes, who are likely juniors or seniors, and who may have already decided whether or not to pursue STEM careers.
@@
Table 3: Shifts in favorable student responses between various questions pairs along with the statistical significance of the shift. Favorable shift is the shift in $p^{\prime}$ values between the questions. Questions #1,3, and 7 deal with student interest in exploring and learning quantum computing while questions #2, 4, and 8 deal with student self-efficacy (confidence) in learning quantum computing.
Question Pairs Favorable Shift Significance of Shift
@@ VI Conclusion
Quantum computing is a critical area of science and technology, and games may play a role in developing the workforce for the field. Our study involving 107 public high school students indicates that playing Qubit Touchdown improved students’ confidence in their ability to learn quantum computing by 20.4%, and learning the game’s connections to quantum computing yielded a net improvement in self-efficacy of 33.4%. Students’ interest in the field, however, was modestly improved by 9.4% by playing the game and learning its technical connections, and we speculate that it may be because the students were mostly juniors and seniors, who may already have plans for after they graduate from high school. Students found the game enjoyable and easy to learn, and found that the game made it easier to learn quantum computing. They recommended it as a learning tool. As such, Qubit Touchdown could be a valuable resource for learning and outreach, giving high students students confidence to pursue technical fields they may have initially considered out of reach.
@@ Acknowledgements.
This material is based upon work supported in part by the National Science Foundation EPSCoR Cooperative Agreement OIA-2044049, Nebraska’s EQUATE collaboration. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.

### v2

# Improving Student Self-Confidence in Quantum Computing with the Qubit Touchdown Board Game

@@
Improving Student Self-Confidence in Quantum Computing with the
@@
Qubit Touchdown is a two-player, competitive board game that was developed to introduce students to quantum computing. A quantum computer is a new kind of computer that is based on the laws of quantum physics, and it can solve certain problems faster than normal computers because it follows a different set of rules. Qubit Touchdown’s game play mirrors the rules of (American) football, with players taking turns moving the football to score the most touchdowns, and no knowledge of quantum computing is needed to play the game. We evaluated the game with 107 public high school students in Precalculus, Advanced Placement (AP) Statistics, and/or AP Physics 1 courses, assessing whether their interest in and self-confidence in their ability to learn quantum computing changed as a result of playing the game and learning about its connections to quantum computing. We also assessed whether the game was easy to learn and enjoyable. We found that students’ interest in quantum computing increased slightly ($p<0.05$), but students’ self-confidence in their ability to learn quantum computing saw greater gains ($p<0.001$); students also widely considered the game accessible and fun. Thus, Qubit Touchdown could be an effective resource to introduce students to Quantum Computing and boost their confidence in learning about the field. Free printables of the game are available, and professionally produced copies can be purchased on demand.
@@ III Connections to Quantum Computing
IV Research on Student Interest, Self-confidence, and Enjoyment
To evaluate whether Qubit Touchdown can affect students’ interest and self-confidence in learning quantum computing, as well as whether the game is easy to learn and fun to play, we conducted a study consisting of 107 public high school students across seven Precalculus, AP Statistics, and/or AP Physics 1 classes over four periods in one school day, which necessitated combining some classes in the library for the activity.
@@ IV Research on Student Interest, Self-confidence, and Enjoyment
Table 2: Overall agree and disagree percentages ($A$ and $D$) for the survey results.
Question # $A$ $D$
Q1 36.4% 26.1%
Q2 36.4% 29.0%
Q3 46.7% 25.2%
Q4 47.7% 15.0%
Q5 84.1% 3.7%
Q6 100.0% 0.0%
Q7 46.7% 22.4%
Q8 69.2% 8.4%
Q9 85.0% 1.9%
Q10 84.1% 3.7%
Q11 78.5% 5.6%
Q12 85.0% 3.7%
Q13 83.2% 1.9%
@@
For a simple first analysis, we treat the Likert scale results as ordinal data and combined agree and strongly agree responses, and disagree and strongly disagree responses, to form overall agree and disagree values, $A$ and $D$, respectively, as shown in Table 2 for all thirteen questions (Q1 to Q13), ignoring neutral responses [21]. The results indicate that students were overall very favorable about their experience of playing Qubit Touchdown. For example, 78.5% of the students reported that playing Qubit Touchdown made quantum computing concepts less intimidating (Q11), and 84.1% of the students reported that they enjoyed playing the game (Q5). Similarly, 83.2% of students recommended that Qubit Touchdown be used as a learning tool for quantum computing (Q13). Students’ interest in exploring quantum computing at different time points (before playing the game, after playing, and after learning the game’s connections to quantum computing) was assessed in Q1, Q3, and Q7, and we compare them by calculating the shift in overall agree percentages, presented as “favorable shift” values (movement toward agreement) in Table 3. After playing the game, students’ interest improved, but learning the game’s connections to quantum computing did not further improve their interest. Similarly, students’ confidence in their ability to learn quantum computing was assessed at the three time points in Q2, Q4, and Q8, and the favorable shifts are also presented in Table 3. Their self-confidence improved after playing the game and again after learning the game’s connection to quantum computing.
@@
Table 3: Shifts in favorable student responses between various questions pairs along with the statistical significance of the shift (Wilcoxon $p$-value). Favorable shift is the shift in $A$ values between the questions ($\Delta A$). Questions 1, 3, and 7 deal with student interest in exploring and learning about quantum computing while questions 2, 4, and 8 deal with student self-confidence in their ability to learn quantum computing.
Question Pairs Favorable Shift ($\Delta A$) Wilcoxon $p$-value
@@
The simple binomial analysis of the Likert scale data above, however, has several disadvantages: it loses the information provided by neutral responses, it loses information from changes in responses from dis/agree to strongly dis/agree or vice-versa, and is not statistically rigorous. Because we have ordinal data which is non-normally distributed and our relevant questions came in sets of three, we first use a Friedman test to determine if the sets of responses to Q1, Q3, and Q7 (as well as Q2, Q4, and Q8) are statistically different when measured at the different time points. We found that there was a significant main effect of measuring student interest in exploring quantum computing at different time points (Q1, Q3, and Q7) by utilizing a Friedman test ($\chi^{2}=8.9589$, $d_{f}=2$, $p<0.05$). In other words, student interest in quantum computing is indeed different before versus after playing the game. In terms of exploring students’ self-confidence in learning quantum computing, we found a stronger effect: there was a significant main effect of time point for confidence in their ability to learn quantum computing by utilizing a Friedman test ($\chi^{2}=49.004$, $d_{f}=2$, $p<0.001$).
Since the Friedman test results were significant and to help control for inflated Type I error due to multiple comparisons, we follow up the Friedman tests with pairwise paired Wilcoxon signed-rank tests with a continuity correction. All $p$-values were two-sided and considered significant if $p<0.05$. The $p$-values from this analysis are reported in Table 3. For students’ interest in quantum computing, there was a statistically significant increase in student interest between Q1 and Q3 ($p<0.05$) and between Q1 and Q7 ($p<0.05$), but not between Q3 and Q7 ($p>0.05$). For students’ self-confidence, there was a statistically significant increase between Q2 and Q4 ($p<0.001$), between Q2 and Q8 ($p<0.001$), and between Q4 and Q8 ($p<0.001$), and they are consistent with the favorable shifts. Table 3 gives both the simple favorable shift (movement towards more agreement) and the Wilcoxon $p$-values for question pairs (which shows that the shift is or is not statistically significant).
In summary, a statistical analysis of the Likert scale data shows that students seem to be both more interested in quantum computing and more self-confident in their ability to learn quantum computing as a result of playing Qubit Touchdown and listening to a short lecture on its connections to quantum computing.
@@ VI Conclusion
Quantum computing is a critical area of science and technology, and games may play a role in developing the workforce for the field. Our study involving 107 public high school students indicates that playing Qubit Touchdown appeared to improve students’ confidence in their ability to learn quantum computing by 11.3%, and learning the game’s connections to quantum computing yielded a net improvement in self-confidence of 32.8%. Students’ interest in the field, however, was modestly improved by 10.3% by playing the game and learning its technical connections, and we speculate that it may be because the students were mostly juniors and seniors, who may already have plans for after they graduate from high school. These shifts in attitudes were found to be statistically significant. Students found the game enjoyable and easy to learn, and found that the game made it easier to learn quantum computing. They recommended it as a learning tool. As such, Qubit Touchdown could be a valuable resource for learning and outreach, giving high students students confidence to pursue technical fields they may have initially considered out of reach.
@@ Acknowledgements.
We thank Dr. Jack Taylor in the Creighton University Biostatistical Core for help with the statistical analysis. This material is based upon work supported in part by the National Science Foundation EPSCoR Cooperative Agreement OIA-2044049, Nebraska’s EQUATE collaboration. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.